# Dynamically localized but deconfined excitations in strained classical spin ice

### v1

@@ III Thermodynamics
We now consider the thermodynamics of our model using classical Monte Carlo simulations. Similar to the case of isotropic spin ice, we find a cross-over from the disordered Coulomb phases to the paramagnet at finite temperature. In contrast, for the symmetry-breaking phases, we identify a critical temperature marking the transition from the paramagnet into the ordered regime. We only conduct simulations at $\delta\geq-1$ ($\eta\geq 0$) due to the duality in Eq. (missing) 4. Our results are summarized in Fig. 4, which contains the specific heat (a-c), monopole density (d-f), and entropy (g,h) for the different regimes discussed previously. Panel (i) shows the magnetic susceptibility measured in the $[111]$ direction. The first column (a,d,g) covers the parameter regime $-1\leq\delta\leq 0$ – the region of the spin ice phase where type 1A monopoles are the lowest-energy excitations. Similarly, the second column (b,e,h) covers $0\leq\delta\leq 1$ – the spin ice phase where type 1B monopoles are lowest in energy. Lastly, the third column (c,f,i) focuses on the ferromagnetic phase, with clear signs of a phase transition occurring at $T_{c}$.
@@
The specific heat, $C_{V}$, contains direct information on the distribution of tetrahedra between the four groups in Fig. 3. Let us start by examining the spin ice phase shown in panels (a,b) of Fig. 4. At infinite temperature, all configurations are equally probable. Upon cooling, higher energy configurations are frozen out. At the level of individual tetrahedra, this means that double monopoles are avoided first, followed by the two types of monopoles. Freezing out these configurations induces a Schottky anomaly, which is visible in the specific heat. When type 1A and 1B monopoles are well separated in energy, each group gives rise to an individual Schottky peak, and as $|\delta|$ increases, we observe the single specific heat peak of $\delta=0$ splitting into two distinct peaks. Note that the entropy released by freezing out these excitations is constant and the heights of the peaks are related to this entropy. As there are more type 1A than type 1B monopoles, the freezing out of type 1A monopoles leads to a larger peak in the specific heat. As $E_{1A}<E_{1B}$ for $\delta<0$, the low-temperature peak in the specific heat is higher in this regime, whereas the opposite is true for $\delta>0$. At $\delta=-1$ ($\delta=1$), type 1A (1B) monopoles are degenerate with the type 0 configurations, and one of the two monopole types is not frozen out, leaving a single peak in the specific heat. When $\delta>1$ we enter the ferromagnetic ground state at low temperatures and observe a sharp peak in the specific heat at the critical temperature $T_{c}$, consistent with a conventional symmetry-breaking phase transition. The critical temperature as a function of $\delta$ can be extracted from the position of the specific heat peak.
@@
The single monopole number density, $n$, is a clear indicator of the phase realized at a certain temperature and distortion $\delta$. Here, we only show the total number of single monopoles, combining types 1A and 1B but not counting the double monopoles. At high temperatures, within the paramagnetic phase, all states are equally likely, yielding $n=8/16=1/2$. In the spin ice phase $-1<\delta<1$, shown in panels (d,e) of Fig. 4, the density drops to zero as ground states are ice configurations without any monopoles. At the phase boundaries, $\delta=\pm 1$, type 0 and type 1B/1A configurations are degenerate, and the density can be well understood by simple state counting. Remarkably, for $\delta=-1$, the darkest red curve in (d), $n$, remains constant at $1/2$. This is due to the degeneracy of the six type 1A and the six type 0 configurations and, in addition, the degeneracy of type 2 and the type 1B configurations. The freezing out of excitations, therefore, leaves no trace in $n$ despite clearly manifesting in the specific heat and entropy. At $\delta=1$, the lightest blue curve in (f), $n$ drops to one quarter as the two type 1B and the six type 0 configurations are degenerate: $n=2/(2+6)=1/4$. In the ferromagnetic regime, for $\delta>1$, the density approaches $n=1$. The monopole density initially drops as the type 1A monopoles are frozen out and then begins to increase at temperatures slightly above $T_{c}$ as the system arranges itself into a type 1B monopole crystal.
@@
Spin ice materials are famously known for their residual entropy. Panels (g,h) in Fig. 4 show the entropy obtained by integrating the specific heat from the high-temperature limit according to
@@
For $|\delta|<1$, $k=6$ and the residual entropy per spin is $s=S/(k_{b}N)=\frac{1}{2}\ln{\left(3/2\right)}$ ). This is also true for $-3<\delta<-1$ since each ice ground state can be mapped to a valid monopole configuration using the transformation in Eq. (missing) 4. At $\delta=1$, the six two-in-two-out configurations have the same energy as the two type 1B monopole configurations. Similarly, at $\delta=-3$, the six type 1A monopole configurations have the same energy as the all-in and all-out (type 2) configurations. Thus, for these two points $k=8$ and $s=\frac{k_{B}}{2}\ln{\left(2\right)}$ ) as illustrated by the lightest blue curve in (h). The residual entropy is even higher at the duality point, $\delta=-1$, with $k=12$ as the six type 0 and six type 1A configurations are all degenerate. This gives the estimate $s=\frac{k_{B}}{2}\ln{\left(3\right)}$ ), consistent with the Monte Carlo result highlighted in dark red in (g).
@@
The susceptibility becomes anisotropic since the distortion in our model breaks the lattice symmetry. Here, we only show the magnetic susceptibility in the $[111]$ direction for $\delta>1$, panel (i) of Fig. 4. It develops sharp peaks at the same temperatures as the specific heat and is a clear sign of a symmetry-breaking phase transition. An analysis of the critical exponents of the susceptibility and specific heat close to $T_{c}$ is provided in App. B.1.
@@ IV Dynamics
The preservation of the extensive ice degeneracy for a broad range of strain in our model, in contrast with previous studies [19, 20], makes it possible to apply the standard model of spin ice dynamics [28, 9]. There, the dynamics of the low-energy excitations is modeled under the assumption that the system evolves through random, independent single-spin flips occurring at a characteristic time scale $\tau_{0}$, cf. App. A for details. More precisely, we perform heat-bath Monte Carlo simulations with uniform single-spin flip attempt rates. This allows us to focus solely on the impact of energetic constraints imposed by strain. Qualitatively similar constraints on monopole motion arise in the “beyond the standard model” for spin ice dynamics [10], albeit from a different origin as dynamical rules on spin flip rates are coming from the quantum mechanical description rather than energetic barriers imposed by strain.
@@ IV.1 Phenomenology
In the following, we focus on time scales smaller than $\kappa_{0}$ and discuss the possible zero-energy moves. Monopole motion is diffusive for $\delta=0$ [10], since the three majority spins are randomly distributed. This situation is in stark contrast to our anisotropic model, where energetic constraints are encoded in the rules governing if a monopole belongs to type 1A or 1B, leading to different levels of restrictions on monopole motion for the two types. Diffusive behavior is recovered in all cases at time scales larger than $\kappa_{0}$, where the distinction between monopole types becomes effectively meaningless.
Type 1B dynamics
@@
Type 1A dynamics
For $\delta<0$, type 1A monopoles are lower in energy. The minority spin of type 1A monopoles is not on sublattice $\mathcal{S}_{[111]}$, making their motion less constrained than that of type 1B monopoles. A majority spin can be flipped without energy cost if the sublattice $\mathcal{S}_{[111]}$ spin in the new tetrahedron hosting the monopole is a majority spin after the flip. In this case, motion in the $[111]$ direction, corresponding to flipping a sublattice $\mathcal{S}_{[111]}$ spin, can always happen at zero energy when there is not already a monopole in the adjacent tetrahedron. Motion through each of the other two majority spins costs zero energy in $2/3$ of cases, cf. Fig. 5. As $2/3$ of the $\mathcal{S}_{K}$ spins are majority spins, the probability that the monopole can move through an arbitrary spin on sublattice $\mathcal{S}_{K}$ is thus $p_{(111)}=4/9$. This is still much below the percolation threshold of the honeycomb lattice. However, monopoles always have the option to propagate perpendicular to the honeycomb layer along $[111]$. This allows them to jump between finite clusters in the $(111)$ plane and leaves the monopoles diffusive with an anisotropic diffusion tensor favoring motion in the $[111]$ direction. In the language of (anisotropic) percolation, we can view their motion as a random walk on a diamond lattice with two filling fractions: $p_{[111]}=1$ and $p_{(111)}=4/9$.
Inverted spin ice
@@ IV.2 Mean-squared displacement
The impact of strain on the monopole motion can be observed by tracking the trajectories of individual monopoles in Monte Carlo simulations. More details are provided in Appendix A. Fig. 6 shows the mean-squared displacement (MSD) of a single, isolated monopole moving without annihilation or creation events. This mimics the natural motion of monopoles at low monopole densities. To capture the anisotropic response of our model, the monopole displacement is projected along the $[111]$ direction, as well as along two orthogonal directions. The choice of these two orthogonal directions is arbitrary and has no impact on the observed MSD. Since creation and annihilation events are banned, the energetics are completely set by the ratio $J\delta T$.
@@
Figure 6: The mean-squared displacement (MSD) of individual isolated monopoles projected along the $[111]$ direction (blue circles) and along two orthogonal directions: $[1\bar{1}0]$ (red down triangles) and $[11\bar{2}]$ (green up triangles). The MSD is measured in units of the squared distance between the center of two neighboring tetrahedra. For $\delta<0$, panel (a), type 1A monopoles are favored, and the monopole dynamics are essentially diffusive, albeit suppressed in the plane normal to $[111]$. The gray lines indicate diffusive behavior, tproportional-$\propto t$, which is isotropic for $\delta=0$ in panel (b). In stark contrast, localization is observed for $\delta>0$, panel (c), where type 1B monopoles are favored. Here, the monopoles are restricted to small clusters on time scales smaller than $\kappa_{0}/\tau_{0}\approx 3\cdot 10^{3}$ (dotted black lines) from Eq. (missing) 7. The localization is followed by diffusive behavior at longer times mediated by the thermal promotion of type 1B monopoles to higher-energy diffusive type 1A monopoles.
@@
Consistent with our phenomenological arguments, anisotropic diffusive monopole motion in the MSD is observed in the type A regime ($\delta\leq 0$; panel (a) of Fig. 6). Motion in the $[111]$ direction is favored as expected. The response in the type B regime ($\delta>0$; panel (c) of Fig. 6) is quite distinct. Here, the MSD initially appears to plateau, indicating that monopoles are localized to small clusters. This is consistent with type 1B monopoles never moving through $\mathcal{S}_{[111]}$ spins, effectively restricting their motion to a small number of tetrahedra within the $(111)$ plane. Indeed, the localization is stronger in the $[111]$ direction [^1] Note that even in the limit of $J|\delta|/T\to\infty$, the MSD projected along $[111]$ will not be zero, as monopole moves through any spins in $\mathcal{S}_{K}$ have a non-zero component along $[111]$.. While type 1B monopoles are initially localized, promotion events occur on a time scale $\kappa_{0}$ and this process restores diffusive monopole motion. In contrast to the type A regime where movement along $[111]$ is favored, the diffusion is isotropic on time scales of $\kappa_{0}$. Once an expensive type 1A monopole is created, its motion is not restricted in any direction because its propagation never enhances the energy. Therefore, as long as no double monopole is created, every move is accepted.
Sub-dimensional diffusion
@@
The classical analog of the ring exchange moves consists of stochastically finding a hexagonal plaquette with aligned spins and flipping the six spins. In our model this lifts the localization of type 1B monopoles, as shown in Fig. 7. The localization is only lifted in the (111) plane as monopoles remain localized in the [111] direction. This is because any ring exchange update, including a $\mathcal{S}_{[111]}$ spin neighboring a monopole, converts the monopole from type 1B to type 1A at an energy cost of $4J\delta$. This process is thermally suppressed for $T\ll 4J\delta$. Thus, when ring exchange updates are included, the monopoles exhibit sub-dimensional diffusion at zero temperature – reminiscent of fractonic motion [14, 13]. We also note that flipping an open string of four spins attached to a type 1B monopole can move it along the $[111]$ direction while preserving the energy.
@@
Figure 7: Left: Mean-squared displacement for a single type 1B monopole in the presence of ring exchange updates that cause zero-energy fluctuation within the ice background. The rate $F$ of ring exchange updates to attempted single-spin flips are given in the legend. Note that this is the rate of performed updates, not the attempt rate. Here, we use moderate ring exchange rates to observe sub-dimensional diffusion at relatively short times. A smaller $F$ leads to sub-dimensional diffusion at late times, and delocalization due to ring exchange would then compete with delocalization due to the thermal promotion to type 1A monopoles. Right: Sub-dimensional diffusion constant plotted versus the ring exchange rate $F$ normalized by the number of spins $N$. The diffusion constant is measured in units of $a_{d}^{2}/\tau_{0}$, where $a_{d}$ is the distance between the center of two neighboring tetrahedra, and was extracted from linear fits to the MSD in the $[1\bar{1}0]$ direction. The sub-dimensional diffusion constant shows expected linear dependence with $F/N$ in the $F\rightarrow 0$ ($\tau_{0}\ll\tau_{lock}$) limit, highlighted in the inset, and the transition to a plateau region at large $F/N$, corresponding to the $\tau_{0}\gg\tau_{lock}$ limit.
@@
The sub-dimensional diffusion of the monopole can be understood in terms of a random walk on a honeycomb lattice where the traversability of a bond changes in time. In a completely unconstrained random walk on a honeycomb lattice, meaning that monopoles can move freely in the (111) plane, diffusion can be observed with a constant $D$: $\langle r^{2}\rangle=Dt$. However, for localized 1B monopoles, moves across any bond are only possible in one of three cases. The additional ring exchange updates scramble the ice background and allow new moves across previously disallowed bonds. The typical time scale, $\tau_{\mathrm{lock}}$, it takes to unlock such a bond scales inversely proportional to the ring exchange rate $F$. Given the single-spin flipping rate $\tau_{0}$, there are two important limits: (i) $\tau_{0}\gg\tau_{\mathrm{lock}}$ and (ii) $\tau_{0}\ll\tau_{\mathrm{lock}}$. For (i), where the ice background is fully scrambled between two attempted hops, the probability of any attempted hop being accepted is $1/3$, and this effectively reduces the diffusion constant to $D/3$. In contrast, for (ii), the time a monopole is localized to a specific cluster is governed by $\tau_{\mathrm{lock}}$. One can model this as monopoles performing a random walk with a new hopping time $\tau_{\mathrm{lock}}$, leading to a reduced diffusion constant $D(\tau_{0}/\tau_{\mathrm{lock}})\propto DF$. In this limit, the diffusion constant thus scales linearly with the ring exchange rate $F$.

### v2

@@ III Thermodynamics
We now consider the thermodynamics of our model using classical Monte Carlo simulations. Similar to the case of isotropic spin ice, we find a cross-over from the disordered Coulomb phases to the paramagnet at finite temperature. In contrast, for the symmetry-breaking phases, we identify a critical temperature marking the transition from the paramagnet into the ordered regime. We only conduct simulations at $\delta\geq-1$ ($\eta\geq 0$) due to the duality in Eq. (missing) 5. Our results are summarized in Fig. 4, which contains the specific heat (a-c), monopole density (d-f), and entropy (g,h) for the different regimes discussed previously. Panel (i) shows the magnetic susceptibility measured in the $[111]$ direction. The first column (a,d,g) covers the parameter regime $-1\leq\delta\leq 0$ – the region of the spin ice phase where type 1A monopoles are the lowest-energy excitations. Similarly, the second column (b,e,h) covers $0\leq\delta\leq 1$ – the spin ice phase where type 1B monopoles are lowest in energy. Lastly, the third column (c,f,i) focuses on the ferromagnetic phase, with clear signs of a phase transition occurring at $T_{c}$.
@@
The specific heat, $C_{V}$, contains direct information on the distribution of tetrahedra between the four groups in Fig. 3. Let us start by examining the spin ice phase shown in panels (a,b) of Fig. 4. At infinite temperature, all configurations are equally probable. Upon cooling, higher energy configurations are frozen out. At the level of individual tetrahedra, this means that double monopoles are avoided first, followed by the two types of monopoles. Freezing out these configurations induces a Schottky anomaly, which is visible in the specific heat. When type 1A and 1B monopoles are well separated in energy, each group gives rise to an individual Schottky peak, and as $|\delta|$ increases, we observe the single specific heat peak of $\delta=0$ splitting into two distinct peaks. Note that the entropy released by freezing out these excitations is constant and the heights of the peaks are related to this entropy. As there are more type 1A than type 1B monopoles, the freezing out of type 1A monopoles leads to a larger peak in the specific heat. As $E_{1A}<E_{1B}$ for $\delta<0$, the low-temperature peak in the specific heat is higher in this regime, whereas the opposite is true for $\delta>0$. At $\delta=-1$ ($\delta=1$), type 1A (1B) monopoles are degenerate with the type 0 configurations, and one of the two monopole types is not frozen out, leaving a single peak in the specific heat.
When $\delta>1$ we enter the ferromagnetic ground state at low temperatures and observe a sharp peak in the specific heat at the critical temperature $T_{c}$, consistent with a conventional symmetry-breaking phase transition. The critical temperature as a function of $\delta$ can be extracted from the position of the specific heat peak. The heat capacity shows more structure in addition to the sharp peak found at the onset of long-ranged ferromagnetic order. Using the $\delta=1.2$ curve as an illustrative example, there are three peaks: a broad low-temperature Schottky peak as type 0 tetrahedra freeze-out, a sharp central peak at $T_{c}$, and another broad Schottky peak at high temperature as type 1A tetrahedra freeze-out. For larger $\delta$ the two Schottky peaks are not independently resolved, but still visible as broadening around the sharp peak. In the thermodynamic limit, we expect the sharp central peak in the specific heat to diverge, while the broad features are system size-independent. See App. C for further discussion.
@@
The single monopole number density, $n$, is a clear indicator of the phase realized at a certain temperature and distortion $\delta$. Here, we only show the total number of single monopoles, combining types 1A and 1B but not counting the double monopoles. At high temperatures, within the paramagnetic phase, all states are equally likely, yielding $n=8/16=1/2$. In the spin ice phase $-1<\delta<1$, shown in panels (d,e) of Fig. 4, the density drops to zero as ground states are ice configurations without any monopoles. At the phase boundaries, $\delta=\pm 1$, type 0 and type 1B/1A configurations are degenerate, and the density can be well understood by simple state counting. Remarkably, for $\delta=-1$, the darkest red curve in (d), $n$, remains constant at $1/2$. This is due to the degeneracy of the six type 1A and the six type 0 configurations and, in addition, the degeneracy of type 2 and the type 1B configurations. The freezing out of excitations, therefore, leaves no trace in $n$ despite clearly manifesting in the specific heat and entropy. At $\delta=1$, the lightest blue curve in (f), $n$ drops to one quarter as the two type 1B and the six type 0 configurations are degenerate: $n=2/(2+6)=1/4$. In the ferromagnetic regime, for $\delta>1$, the density approaches $n=1$. The total monopole density initially drops as the type 1A configurations are converted into type 0 and type 1B configurations. The density then begins to increase at temperatures slightly above $T_{c}$ as the system arranges itself into a type 1B monopole crystal.
@@
Spin ice, in all its incarnations, is famously known for its residual entropy. Panels (g,h) in Fig. 4 show the entropy obtained by integrating the specific heat from the high-temperature limit according to
@@
For $|\delta|<1$, $k=6$ and the residual entropy per spin is $s=S/(k_{b}N)=\frac{1}{2}\ln{\left(3/2\right)}$ ). This is also true for $-3<\delta<-1$ since each ice ground state can be mapped to a valid monopole configuration using the transformation in Eq. (missing) 5. At $\delta=1$, the six two-in-two-out configurations have the same energy as the two type 1B monopole configurations. Similarly, at $\delta=-3$, the six type 1A monopole configurations have the same energy as the all-in and all-out (type 2) configurations. Thus, for these two points $k=8$ and $s=\frac{k_{B}}{2}\ln{\left(2\right)}$ ) as illustrated by the lightest blue curve in (h). The residual entropy is even higher at the duality point, $\delta=-1$, with $k=12$ as the six type 0 and six type 1A configurations are all degenerate. This gives the estimate $s=\frac{k_{B}}{2}\ln{\left(3\right)}$ ), consistent with the Monte Carlo result highlighted in dark red in (g).
@@
The susceptibility becomes anisotropic since the distortion in our model breaks the lattice symmetry. Here, we only show the magnetic susceptibility in the $[111]$ direction for $\delta>1$, panel (i) of Fig. 4. It develops sharp peaks at the same temperatures as the specific heat and is a clear sign of a symmetry-breaking phase transition. An analysis of the critical exponents of the susceptibility and specific heat close to $T_{c}$ is provided in App. C.1. We also plot the susceptibility at $\delta=1$ (light blue) for comparison, which increases as $\propto T^{-1}$ upon cooling.
@@ IV Dynamics
The preservation of the extensive ice degeneracy for a broad range of strain in our model, in contrast with previous studies [19, 20], where strain along the $[100]$ crystallographic axis was considered, makes it possible to apply the standard model of spin ice dynamics [29, 9]. There, the dynamics of the low-energy excitations is modeled under the assumption that the system evolves through random, independent single-spin flips occurring at a characteristic timescale $\tau_{0}$, cf. App. B for details. More precisely, we perform heat-bath Monte Carlo simulations with uniform single-spin flip attempt rates. This is the simplest form of dynamics one can imagine, and therefore allows us to focus solely on the impact of energetic constraints imposed by strain, as opposed to any dynamical constraints imposed by spin flipping rules. Qualitatively similar constraints on monopole motion arise in the “beyond the standard model” for spin ice dynamics [10], albeit from a different origin as dynamical rules on spin flip rates come from the quantum mechanical description rather than energetic barriers imposed by strain. The beyond the standard model is commented on further in Sec. V, but will not be used in this work.
@@ IV.1 Phenomenology
In the following, we focus on timescales smaller than $\kappa_{0}$ and discuss the possible zero-energy moves. Monopole motion is diffusive for $\delta=0$ [10], since the three majority spins are randomly distributed. This situation is in stark contrast to our anisotropic model, where energetic constraints are encoded in the rules governing if a monopole belongs to type 1A or 1B, leading to different levels of restrictions on monopole motion for the two types. Diffusive behavior is recovered in all cases at timescales larger than $\kappa_{0}$, as this is the typical timescale on which monopoles change type. On this timescale lower-energy monopoles thus overcome energy barriers associated with changing type, which opens up further paths for them resulting in diffusive monopole motion.
Type 1B dynamics.
@@
Type 1A dynamics.
For $\delta<0$, type 1A monopoles are lower in energy. The minority spin of type 1A monopoles is not on sublattice $\mathcal{S}_{[111]}$, making their motion less constrained than that of type 1B monopoles. A majority spin can be flipped without energy cost if the sublattice $\mathcal{S}_{[111]}$ spin in the new tetrahedron hosting the monopole is a majority spin after the flip. In this case, motion through a sublattice $\mathcal{S}_{[111]}$ spin, can always happen at zero energy when there is not already a monopole in the adjacent tetrahedron. Motion through each of the other two majority spins costs zero energy in $2/3$ of cases, cf. Fig. 5. As $2/3$ of the $\mathcal{S}_{k}$ spins are majority spins, the probability that the monopole can move through an arbitrary spin on sublattice $\mathcal{S}_{k}$ is thus $p_{(111)}=4/9$. This is still much below the percolation threshold of the honeycomb lattice. However, monopoles can propagate perpendicular to the honeycomb layer through the $\mathcal{S}_{[111]}$ spins. This allows them to jump between finite clusters in the $(111)$ plane and leaves the monopoles diffusive with an anisotropic diffusion tensor favoring motion in the $[111]$ direction. In the language of (anisotropic) bond percolation, we can view their motion as a random walk on a diamond lattice with two filling fractions: $p_{[111]}=1$ for bonds along the $[111]$ direction and $p_{(111)}=4/9$ for the three remaining bond directions.
We note that monopoles can, in principle, become dynamically localized in this regime. This occurs when all six moves taking a monopole away from two tetrahedra connected through the $\mathcal{S}_{[111]}$ spin are blocked. The monopole is then only able to move between these two tetrahedra. Instances like these are rare, and we find that the vast majority of type 1A monopoles are not dynamically localized.
Inverted spin ice.
@@ IV.2 Mean-squared displacement
The impact of strain on the monopole motion can be observed by tracking the trajectories of individual monopoles in Monte Carlo simulations. More details are provided in Appendix B. Fig. 6 shows the mean-squared displacement (MSD) of a single, isolated monopole moving without annihilation or creation events. This mimics the natural motion of monopoles at low monopole densities. To capture the anisotropic response of our model, the monopole displacement is projected along the $[111]$ direction, as well as along two orthogonal directions. The choice of these two orthogonal directions is arbitrary and has no impact on the observed MSD. Since creation and annihilation events are banned, the energetics are completely set by the ratio $J\delta/T$.
@@
Figure 6: The mean-squared displacement (MSD) of individual isolated monopoles projected along the $[111]$ direction (blue circles) and along two orthogonal directions: $[1\bar{1}0]$ (red down triangles) and $[11\bar{2}]$ (green up triangles). The MSD is measured in units of the squared distance between the center of two neighboring tetrahedra. For $\delta<0$, panel (a), type 1A monopoles are favored, and the monopole dynamics are essentially diffusive, albeit suppressed in the plane normal to $[111]$. The gray lines indicate diffusive behavior, tproportional-$\propto t$, which is isotropic for $\delta=0$ in panel (b). In stark contrast, localization is observed for $\delta>0$, panel (c), where type 1B monopoles are favored. Here, the monopoles are restricted to small clusters on timescales smaller than $\kappa_{0}/\tau_{0}\approx 3\cdot 10^{3}$ (dotted black lines) from Eq. (missing) 8. The localization is followed by diffusive behavior at longer times mediated by the thermal promotion of type 1B monopoles to higher-energy diffusive type 1A monopoles.
@@
Consistent with our phenomenological arguments, anisotropic diffusive monopole motion in the MSD is observed in the type A regime ($\delta\leq 0$; panel (a) of Fig. 6). Motion in the $[111]$ direction is favored as expected. The response in the type B regime ($\delta>0$; panel (c) of Fig. 6) is quite distinct. Here, the MSD initially appears to plateau, indicating that monopoles are localized to small clusters. This is consistent with type 1B monopoles never moving through $\mathcal{S}_{[111]}$ spins, effectively restricting their motion to a small number of tetrahedra within the $(111)$ plane. Indeed, the localization is stronger in the $[111]$ direction [^1] Note that even in the limit of $J|\delta|/T\to\infty$, the MSD projected along $[111]$ will not be zero, as monopole moves through any spins in $\mathcal{S}_{k}$ have a nonzero component along $[111]$.. While type 1B monopoles are initially localized, promotion events occur on a timescale $\kappa_{0}$, and this process restores diffusive monopole motion. In contrast to the type A regime where movement along $[111]$ is favored, the diffusion is isotropic on timescales of $\kappa_{0}$. Once an expensive type 1A monopole is created, its motion is not restricted in any direction because its propagation never enhances the energy. Therefore, as long as no double monopole is created, every move is accepted.
Subdimensional diffusion.
@@
The classical analog of the ring exchange moves consists of stochastically finding a hexagonal plaquette with aligned spins and flipping the six spins. In our model, this lifts the localization of type 1B monopoles, as shown in Fig. 7. The localization is only lifted in the (111) plane as monopoles remain localized in the [111] direction. This is because any ring exchange update, including a $\mathcal{S}_{[111]}$ spin neighboring a monopole, converts the monopole from type 1B to type 1A at an energy cost of $4J\delta$. This process is thermally suppressed for $T\ll 4J\delta$. Thus, when ring exchange updates are included, the monopoles exhibit subdimensional diffusion at zero temperature – reminiscent of fractonic motion [14, 13]. We also note that flipping an open string of four spins attached to a type 1B monopole can move it along the $[111]$ direction while preserving the energy.
@@
Figure 7: Left: Mean-squared displacement for a single type 1B monopole in the presence of ring exchange updates that cause zero-energy fluctuation within the ice background. The rate $F$ of ring exchange updates to attempted single-spin flips are given in the legend. Note that this is the rate of performed updates, not the attempt rate. Here, we use moderate ring exchange rates to observe subdimensional diffusion at relatively short times. A smaller $F$ leads to subdimensional diffusion at late times, and delocalization due to ring exchange would then compete with delocalization due to the thermal promotion to type 1A monopoles. Right: Subdimensional diffusion constant plotted versus the ring exchange rate $F$ normalized by the number of spins $N$. The diffusion constant is measured in units of $a_{d}^{2}/\tau_{0}$, where $a_{d}$ is the distance between the center of two neighboring tetrahedra, and was extracted from linear fits to the MSD in the $[1\bar{1}0]$ direction. The subdimensional diffusion constant shows expected linear dependence with $F/N$ in the $F\rightarrow 0$ ($\tau_{0}\ll\tau_{lock}$) limit, highlighted in the inset, and the transition to a plateau region at large $F/N$, corresponding to the $\tau_{0}\gg\tau_{lock}$ limit.
@@
The subdimensional diffusion of the monopole can be understood in terms of a random walk on a honeycomb lattice where the traversability of a bond changes in time. In a completely unconstrained random walk on a honeycomb lattice, meaning that monopoles can move freely in the (111) plane, diffusion can be observed with a constant $D$: $\langle r^{2}\rangle=Dt$. However, for localized 1B monopoles, moves across any bond are only possible in one of three cases. The additional ring exchange updates scramble the ice background and allow new moves across previously disallowed bonds. The typical timescale, $\tau_{\mathrm{lock}}$, it takes to unlock such a bond scales inversely proportional to the ring exchange rate $F$. Given the single-spin flipping rate $\tau_{0}$, there are two important limits: (i) $\tau_{0}\gg\tau_{\mathrm{lock}}$ and (ii) $\tau_{0}\ll\tau_{\mathrm{lock}}$. For limit (i), where the ice background is fully scrambled between two attempted hops, the probability of any attempted hop being accepted is $1/3$, and this effectively reduces the diffusion constant to $D/3$. In contrast, for limit (ii), the time a monopole is localized to a specific cluster is governed by $\tau_{\mathrm{lock}}$. One can model this as monopoles performing a random walk with a new hopping time $\tau_{\mathrm{lock}}$, leading to a reduced diffusion constant $D(\tau_{0}/\tau_{\mathrm{lock}})\propto DF$. In this limit, the diffusion constant thus scales linearly with the ring exchange rate $F$.